Use an appropriate trigonometry substitution to evaluate ∫▒√(x^2+16)/x^4  dx


1
Expert's answer
2021-05-07T14:14:27-0400

We will calculate "\int(\sqrt(x^2+16))dx/(x^4)"

put x=4tan(t)

So dx="4sec^2tdt"

Now integration becomes "\int((4\sqrt(1+tan^2t))*4sec^2t)dt/(4^4tan^4(t))"

="\int(sec^3(t) dt)/(16tan^4(t))"

="\int(1/cos^3(t))*(16(cos^4(t)/sin^4(t)))"

="\int(cos(t))/(16sin^4(t))"

Now let sin(t)=z So cos(t)dt=dz

So integration becomes "\int(dz/16z^4) =\int((z^(-4)dz))/16"

= -1/(48z^3)

=-1/(48sin^3(t)) As "\int(x^ndx)=(x^(n+1))/(n+1)" We had taken x=4tan(t) ,so tan(t)=x/4 ,in a right angle triangle tan(t)=p/b , so p=x & b=4

So h="\sqrt(x^2+16)" ,So "sin(t)=p/h=x/(\sqrt (x^2+16))" "=-((\sqrt(x^2+16))^3)/(48x^3) + C" (ans)

Here C is constant of integration.